# Oblique photons, plasmons, and current-plasmons in relativistic plasmas and their topological implications

### v1

@@ Abstract
Photons in vacuum are transverse in any inertial frame; longitudinal photons only exist virtually. By developing a manifestly covariant theory for electromagnetic excitations in relativistic plasmas and applying Wigner’s little group method for elementary particle classifications, we show that photons in plasmas are neither transverse nor longitudinal; they are oblique. Plasmons are electromagnetic and oblique as well. The Lorentz invariant characteristics that distinguishes photons and plasmons is covariant compressibility. The manifestly covariant theory predicts the existence of the current-plasmon, a third oblique, electromagnetic eigenmode, and it also enables the study of photon topology in plasmas. Plasmas remove the photon’s Dirac point in vacuum by giving it an effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon.
@@
Previous work has considered the electromagnetic waves and electrostatic waves in relativistic plasmas [1, 2, 3, 4, 5, 6]. However, despite our best effort, we have found no systematic, manifestly covariant treatment of the subject in the existing literature. Most previous studies focused on relativistic corrections or effects on specific modes, and dispersion relations of the modes were often written down by heuristic arguments based on the requirement of covariance or the necessity of the relativistic factor $\gamma$ and with a priori assumptions about the dispersion relations and polarization properties. Because of the lacking of a systematic, manifestly covariant treatment, certain important phenomena, such as the self-induced transparency, have been debated about in the literature [1, 2, 7, 8, 3, 4, 5, 9, 6]. With the arrival of increasingly powerful laser facilities, we hope the results reported in the present study contribute to the understanding on this subject.
@@
Using the manifestly covariant theory and classification techniques developed, we rigorously show that the system admits one 2D mode that can be identified as a photon and one 1D mode that can be identified as a plasmon, as well as their anti-particles. But the photons are neither transverse nor longitudinal. The electric field of photons is oblique to its momentum. Plasmons are not electrostatic, and they are also oblique. In addition, the manifestly covariant theory also predicts the existence of the current-plasmon, which is a third, previously unknown oblique electromagnetic excitation in relativistic plasmas. If photons, plasmons, and current-plasmons all are electromagnetic and oblique in general inertial frames, is there a covariant label that can be used to distinguish them? We find that covariant compressibility $\chi$ is such a label. Photons and current-plasmons are not covariantly compressible, but plasmons are. The manifestly covariant theory also enables the study of photon topology in plasmas. We will show that plasmas remove the photon’s upright Dirac point in vacuum by giving it an effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon.
@@
    which is expected because $\omega^{2}-k_{j}k^{j}$ is a Lorentz invariant scalar. The dispersion relations look identical to their counterparts, Eq. (42), in the rest frame $x^{\prime\prime}.$ However, the polarization properties of the eigenmodes change under the Lorentz transformations. In the frame $x^{\prime},$ photons are still transverse because $\Lambda^{\prime-1}\in\text{SO}(3),$ which rotates both $\delta\boldsymbol{E}^{\prime\prime}$ and $\boldsymbol{k}^{\prime\prime}$as 3D vectors and preserves the transversality. On the other hand, $\Lambda^{-1}$ is a boost, and it transforms $\boldsymbol{k}^{\prime}$ as the spatial components of the contravariant 4-vector $k^{\prime}$ and $\delta\boldsymbol{E}^{\prime}$ as the temporal components of the electromagnetic tensor $F$. The transversality is broken by a general boost $\Lambda^{-1}$ unless $\omega_{p}$ vanishes. To see this, let’s take the example of

### v2

@@ Abstract
Photons in vacuum are transverse in any inertial frame; longitudinal photons only exist virtually. By developing a manifestly covariant theory for electromagnetic excitations in relativistic plasmas and applying Wigner’s little group method for elementary particle classifications, we show that photons in plasmas are neither transverse nor longitudinal; they are oblique. Plasmons are electromagnetic and oblique as well. The Lorentz invariant characteristics that distinguishes photons and plasmons is covariant compressibility. The manifestly covariant theory predicts the existence of the current-plasmon, a third oblique, electromagnetic eigenmode, and it also enables the study of photon topology in plasmas. Plasmas remove the photon’s Dirac point in vacuum by giving it an effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon. The manifest covariance of the theory demonstrates that relativistic transparency, despite being widely studied, does not exist in plasmas.
@@
Previous work has considered the electromagnetic waves and electrostatic waves in relativistic plasmas [1, 2, 3, 4, 5, 6]. However, despite our best effort, we have found no systematic, manifestly covariant treatment of the subject in the existing literature. Most previous studies focused on relativistic corrections or effects on specific modes, and dispersion relations of the modes were often written down by heuristic arguments based on the requirement of covariance or the necessity of the relativistic factor $\gamma$ and with a priori assumptions about the dispersion relations and polarization properties. Because of the lacking of a systematic, manifestly covariant treatment, certain important phenomena, such as the self-induced transparency in relativistic plasmas, have been debated about in the literature [1, 2, 7, 8, 3, 4, 5, 9, 6]. With the arrival of increasingly powerful laser facilities, we hope the results reported in the present study contribute to the understanding on this subject.
@@
Using the manifestly covariant theory and classification techniques developed, we rigorously show that the system admits one 2D mode that can be identified as a photon and one 1D mode that can be identified as a plasmon, as well as their anti-particles. But the photons are neither transverse nor longitudinal. The electric field of photons is oblique to its momentum. Plasmons are not electrostatic, and they are also oblique. In addition, the manifestly covariant theory also predicts the existence of the current-plasmon, which is a third, previously unknown oblique electromagnetic excitation in relativistic plasmas. If photons, plasmons, and current-plasmons all are electromagnetic and oblique in general inertial frames, is there a covariant label that can be used to distinguish them? We find that covariant compressibility $\chi$ is such a label. Photons and current-plasmons are not covariantly compressible, but plasmons are. The manifestly covariant theory also enables the study of photon topology in plasmas. We will show that plasmas remove the photon’s upright Dirac point in vacuum by giving it an effective mass, but create a tilted Dirac-Weyl point by reviving the virtual longitudinal photon. In addition, the manifest covariance of the theory demonstrates that relativistic transparency [1, 2, 7, 8, 3, 4, 5, 9, 6], despite being widely studied, does not exist in plasmas.
@@
    which is expected because $\omega^{2}-k_{j}k^{j}$ is a Lorentz invariant scalar. The dispersion relations are identical to their counterparts, Eq. (42), in the rest frame $x^{\prime\prime}.$ A direct consequence of this result is that the cutoff frequency for photons in a relativistic plasma does not depend on how fast the plasma is moving in the lab frame $x$. The manifest covariance of the dispersion relation demonstrates that relativistic transparency [1, 2, 7, 8, 3, 4, 5, 9, 6], despite being widely studied, does not exist in plasmas. However, the polarization properties of the eigenmodes change under the Lorentz transformations. In the frame $x^{\prime},$ photons are still transverse because $\Lambda^{\prime-1}\in\text{SO}(3),$ which rotates both $\delta\boldsymbol{E}^{\prime\prime}$ and $\boldsymbol{k}^{\prime\prime}$as 3D vectors and preserves the transversality. On the other hand, $\Lambda^{-1}$ is a boost, and it transforms $\boldsymbol{k}^{\prime}$ as the spatial components of the contravariant 4-vector $k^{\prime}$ and $\delta\boldsymbol{E}^{\prime}$ as the temporal components of the electromagnetic tensor $F$. The transversality is broken by a general boost $\Lambda^{-1}$ unless $\omega_{p}$ vanishes. To see this, let’s take the example of